Write the equation -2x + 6y = 9 in polar form.
9sqrt(10)/20 = r cos (theta - 108)
9sqrt(10)/20 = r cos (theta - 72)
7sqrt(10)/20 = r cos (theta - 108)
7sqrt(10)/20 = r cos (theta - 72)
I'm really stuck here so any help would be very much appreciated, thanks!

Answer by rothauserc(4718)

-2x +6y = 9
:
standard form is
:
-2x +6y -9 = 0 then
2x -6y +9 = 0
:
A=2, B=-6, C=9
:
+ or - square root(4+36) = + or - 2square root(10)
:
use -2square root(10), the minus is opposite sign of C
:
(2x /(-2square root(10)) -6y/(-2square root(10)) +(9 / (-2square root(10)) = 0
:
multiply the numerator and denominator of each fraction by square root(10)
:
(-2square root(10) / 20)x +(6square root(10) / 20)y -(9square root(10) / 20) = 0
:
note that -cos, +sin implies quadrant II
:
phi = cos^(-1) of (-2square root(10) / 20) = 108.434948822
:
**********************************************************
correct answer is
(9 * square root(10) / 20) = r * cos(theta - 108)
:
note that * means multiply
*********************************************************